During the first 3 weeks 0 the televis on season, the Saturday evening 8:00 PM. t 9:00 n au lence pro o ons ere o ded as ABC 30% С 5 27 o NBC 2 o and independents 2 revision yielded uie allowing viewing dudience daLa: A C homes, CBS 67 homes, NBC 8 hornes arid independents 57 h rne. Test with α to de ermine wheth A sample of homes t eeks after a Saturday n gnt schedule er the viewing audience pr oportions didnued. Round your answers to two decimal places Test statistic p-value is betweernSelect your answer-v Select your answer
Answer #1

Test statistic is

(O - E)2

O: Observed frequency
E: Expected frequency.

E = n*pi

n = 300

O p E (O-E) (O-E)^2 (O-E)^2/E
95 0.3 90 5 25 0.277778
67 0.27 81 -14 196 2.419753
81 0.22 66 15 225 3.409091
57 0.21 63 -6 36 0.571429
Total 300 6.68

\chi ^{2}=6.68

Degrees of freedoms = Number of E's - 1 = 4 - 1 = 3

P-value is between 0.05 and 0.1

P-value > 0.05 we fail to reject null hypothesis.

Conclusion: The viewing audience proportions not changed.